OpenStudy (anonymous):

the square root of x^(16) is x^(4)

OpenStudy (amistre64):

only if 16/2 = 4

OpenStudy (anonymous):

can you show me how u get the answer

OpenStudy (amistre64):

i just did ....

OpenStudy (amistre64):

\[\Large\sqrt{x^n}=x^{\frac n2}\]

OpenStudy (amistre64):

youve got n=16 in this case